# On-the-fly Surrogation for Complex Nonlinear Dynamics*

### v2

@@ III-B Interpolation of local system dynamics
As mentioned previously, in our problem setting, we only have a finite number of local snapshots of the system. Specifically, we aim to leverage the fact that the system linearizations contained in $\mathcal{D}$ according to (4) provide an exact description of the original system (1) at the linearization points. Based on this reasoning, solving the surrogation problem amounts to finding a close approximation of $\mathcal{L}_{f}$ using the set $\mathcal{D}$ while introducing no error at the linearization points. This requirement is formulated as an interpolation problem as follows.
@@ III-B1 Formulation of the interpolation problem
satisfies the following conditions
$r(z)$ $=0\quad\text{for all}\quad z^{(i)}\in\Theta;$ (14a)
@@
The problem of determining $\mathcal{I}^{\mathcal{\mathcal{D}}}$ subject to the requirements (14) is a multidimensional interpolation problem suggesting that regression-based methods, such as [19, 20, 21], are not suitable since they do not satisfy (14a).
Furthermore, one must not assume that the observation points contained in $\mathcal{D}$ are uniformly distributed. In reality, they are scattered instead.
@@ III-B2 Multidimensional scattered interpolation with RBFs
Although there exist multiple alternatives to find $\mathcal{I}^{\mathcal{\mathcal{D}}}$ based on the given multidimensional scattered set $\mathcal{D}$, we propose the use of RBFs for this purpose, as this methodology has been proven to be an efficient and robust solution of the multidimensional scattered interpolation problem [22, 23, 24, 25]. The RBF-based interpolant that approximates $\mathcal{L}_{f}$ is given by
@@
where $\phi$ is a RBF, i.e., a continuous function $\phi:[0,\infty)\to\mathbb{R}$ which is invariant under translation and rotation, $\alpha\in\mathbb{R}^{N\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$, and the interpolation condition in (14a) is imposed as
@@
where $\mathbb{I}_{a}^{b}\coloneq\{i\in\mathbb{Z}^{+}\mid a\leq i\leq b\}$ is an index set. Then, the coefficients $\alpha_{i}$ can be determined by formulating (16) as the following equation system:
$R\;\alpha=\gamma,$ (17a)
with
$[R]_{i,j}$ :-$\coloneq\phi(z^{(i)}-z^{(j)}),:$ $i,j\in\mathbb{I}_{1}^{N};$ (17b)
$[\gamma]_{i}$ :-$\coloneq\operatorname{vec}(M_{i}),:$ $i\in\mathbb{I}_{1}^{N},$ (17c)
where $R\in\mathbb{R}^{N\times N}$ and $\gamma\in\mathbb{R}^{N\times{n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}}$.
@@ III-B3 Extension of the RBF interpolant with a polynomial tail
When variations of the local system dynamics are described by polynomial relationships of total degree less than $m$ the RBF-based interpolant (15) can be improved by adding a polynomial tail, as introduced in [14]. This polynomial tail represents the basis of $d$-variate polynomials of total degree at most $m-1$ which guarantees polynomial precision. Mathematically, the polynomial tail is denoted by $\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}$ and is defined as
$\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z)=\operatorname{mat}_{n_{\mathrm{x}}}(\sum_{j=1}^{Q}\beta_{j}q_{j}(z)),$ (18)
where $\beta\in\mathbb{R}^{Q\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ and $q_{1},\dots,q_{Q}$ is a basis for the $d$-variate polynomials of degree $\leq m-1$ with dimension
$Q=\frac{(m-1+d)!}{(m-1)!\,d!},\quad\text{with}\quad d=n_{\mathrm{x}}+n_{\mathrm{u}}+n_{\mathrm{\eta}}.$ (19)
In this case, the interpolant that approximates $\mathcal{L}_{f}$ becomes
$\mathcal{I}^{\mathcal{D}}(z)=\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}(z)+\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z),$ (20)
@@
$\mathcal{I}^{\mathcal{D}}(z^{(i)})=M_{i},\quad i\in\mathbb{I}_{1}^{N},\$ (21)
@@ III-B4 Determining coefficients of the interpolant
$\begin{bmatrix}R&P\\ P^{\top}&0\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\begin{bmatrix}\gamma\\ 0\end{bmatrix},$ (23a)
@@
$[P]_{i,j}$ :-$\coloneq q_{j}(z^{(i)}),\quad i\in\mathbb{I}_{1}^{N},\ j\in\mathbb{I}_{1}^{Q},:$ (23b)
where $P\in\mathbb{R}^{N\times Q}$, and $R$ and $\gamma$ are defined as in (17).
Now, we can show the uniqueness of solutions of (23) on the basis of [14, Chapter 8]. To this end, we first introduce the notion of conditionally positive definite (CPD) functions [14, Definition 8.1].
@@
Then, using Definition 3, we have the following result that closely follows the treatment of [14, Theorem 8.21].
@@
Suppose that the function $\phi$ is a CPD of order $m$ and the columns of $P$ are linearly independent, where $P$ is given in (23b). Then, (23) has a unique solution.
@@
Suppose that $[\alpha^{\top}\ \beta^{\top}]^{\top}\in\mathbb{R}^{(N+Q)\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ is a solution of the homogeneous equation system in (23), i.e., with $\gamma=0$. Then, we have
$R\alpha+P\beta$ $=0;$ (26)
@@
where the second equation means that $\alpha$ satisfies the side condition (22). Now, pre-multiplying the first equation by $\alpha^{\top}$ gives $\alpha^{\top}R\alpha+\alpha^{\top}P\beta=\alpha^{\top}R\alpha+\left(P^{\top}\alpha\right)^{\top}\beta=\alpha^{\top}R\alpha=0$. As $\phi$ is CPD of order $m$, we can conclude that $\alpha=0$ and thus $P\beta=0$. Then, since the columns of $P$ are linearly independent, we can conclude that $\beta=0$. Therefore, (26) has trivial solutions only, meaning that the left-hand-side matrix of (23a) is full column rank, and thus (23) has a unique solution. ∎
Based on the results in Theorem 4, if the chosen RBF is CPD and the data points $z^{(i)}$ are assumed to be collected in a manner such that P in (23b) is full rank, (23) has a unique solution, which guarantees that, for a given $\mathcal{D}$, there exists a unique RBF interpolant extended with a unique polynomial tail in (20) that satisfies the interpolation condition (14a). One example of a CPD RBF function is the well-known Hardy multiquadric [14], which is CPD of (minimum) order $m=1$, given by
$\phi(z,c)=-\left(c^{2}+\|z\|_{2}^{2}\right)^{1/2},$ (27)
@@
Now, for a given data set $\mathcal{D}$, we can use (15) to approximate the linearization operator $\mathcal{L}_{f}$ within the admissible space. Additionally, if the nonlinear system (1) is expected to contain polynomial relationships in its dynamics, we can use (20) instead to guarantee polynomial precision. In either case, the construction of the interpolant $\mathcal{I}^{\mathcal{D}}$ only involves an unconstrained linear system of equations, which can be efficiently solved with standard linear least squares methods [26]. In the next section, we detail how to utilize the proposed interpolant $\mathcal{I}^{\mathcal{D}}$ in combination with the results from Theorem 1 to achieve our objective of constructing our surrogate model on-the-fly .
@@ III-C Finding the surrogate representation on-the-fly
$\hat{\Sigma}\coloneq\bigl{\{}\xi\hat{x}=\hat{f}(\hat{x},u,\eta)=\hat{F}(\hat{x},u,\eta)\begin{bmatrix}\hat{x}\\ u\end{bmatrix},$ (28a)
@@
[…]
and the integral in (28b) is left unevaluated. Lastly, this representation can be used to assemble the surrogate model on-the-fly, i.e., the surrogate model is only computed when it is required by solving the definite integral in (28b), as summarized in Algorithm 2. For this, it is necessary to invoke a numerical integration scheme to solve (28b) in a computationally efficient manner, as an analytic solution can be costly and often not computable.
@@
Input: Data set $\mathcal{D}$ of the form (4), RBF function $\phi(z,c)$, width parameter $c$, polynomial basis $q(z)$ of total degree $m$.
@@
2:Find the coefficients $\alpha$ and $\beta$ by solving the system in (23).
@@
1:Formulate the surrogate representation $\hat{f}$ in (28a).
2:Compute $\xi\hat{x}(t)=\hat{f}(\hat{x}(t),u(t),\eta)$, solving the integral in (28b) with the numerical integration scheme.
@@
The assembly of the surrogate model on-the-fly is a key feature of the proposed approach, as the inclusion of new data points only requires the computation of $\mathcal{I}^{\mathcal{D}}$ and thus is straightforward. This property is particularly advantageous for updating the surrogate model upon the acquisition of new system observations, which is a common scenario in real-world applications. In addition, the proposed approach guarantees that $\hat{f}$ converges to the true nonlinear system $f$ as the number of observations in $\mathcal{D}$ increases, leading to the following result.
@@
For a nonlinear system $\Sigma$ given by (1) where $f\in\mathcal{C}^{k}$ with $k\in\mathcal{Z}^{+}$, a set $\mathcal{D}=\{\Theta,\mathcal{M}\}$ given by (4) with $N\in\mathbb{Z}^{+}$ distinct linearizations, the interpolant $\mathcal{I}^{\mathcal{D}}$ given by (20) that approximates $\mathcal{L}_{f}$ in (2), and the surrogate $\hat{f}$ given by (28a), for a given point $(z_{\ast})\in\mathcal{Z}$, it holds that
@@
as $N\to\infty$ with all linearization points considered to be distinct.
@@
The proof follows from the fact that $\mathcal{I}^{\mathcal{D}}$ satisfies the interpolation condition in (14a), therefore $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|\neq 0$ at any point $z\in\mathcal{Z}_{\Theta}^{\mathrm{c}}$, where $\mathcal{Z}_{\Theta}^{\mathrm{c}}$ denotes the relative complement of $\mathcal{Z}$ w.r.t. $\Theta$, i.e., $\mathcal{Z}_{\Theta}^{\mathrm{c}}\coloneq\mathcal{Z}\setminus\Theta$. Then, as $N\to\infty$, #$\#\mathcal{Z}_{\Theta}^{\mathrm{c}}\to 0$, where ##\## denotes the cardinality of the set. Thus, $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|\to 0$, and consequently, $\|f(z_{\ast})-\hat{f}(z_{\ast})\|\to 0$ by Theorem 1. ∎
@@ III-D Optimization of the RBFs width parameter
In most cases, the RBFs are dependent on a width parameter $c$, for instance, see (27). The value of $c$ plays a major role in the achievable approximation performance of the interpolant $\mathcal{I}^{\mathcal{D}}$ in between observation points, and consequently, the accuracy of the surrogate $\hat{f}$. Therefore, based on the “leave-one-out” cross validation approach proposed in [15], we formulate the following optimization procedure to find $c$. For a given set of linearizations $\mathcal{D}=\{z^{(k)},M_{k}\}$, an RBF $\phi(z,c)$ and a polynomial basis $q$ of total degree $m$,
$c^{\ast}=\min_{c}\|[E_{1}\ldots E_{N}](c)\|_{p},$ (30)
@@
where $I^{\mathcal{D}_{k}}$ denotes the interpolant given by (20) over the set $\mathcal{D}_{k}\coloneq\mathcal{D}\setminus\{z^{(k)}\}$, and $\|\cdot\|_{p}$ denotes the matrix induced $p$-norm. In particular, $p=\infty$ is in line with the objective (14b).

### v3

@@ III-B Interpolation of local system dynamics
The system linearizations in $\mathcal{D}$ (4) provide exact representations of the original system (1) at their respective linearization points. The surrogation problem then amounts to approximating $\mathcal{L}_{f}$ from $\mathcal{D}$ while preserving this exactness, which naturally leads to the following interpolation problem.
@@ III-B1 Formulation of the interpolation problem
satisfies the following conditions:
$r(z)$ $=0\quad\text{for all}\quad z^{(i)}\in\Theta,$ (14a)
@@
The problem of determining $\mathcal{I}^{\mathcal{\mathcal{D}}}$ subject to (14) is a multidimensional interpolation problem suggesting that most regression-based methods, such as [17, 18], are not suitable since they do not satisfy (14a). Furthermore, one must not assume that the observation points contained in $\mathcal{D}$ are uniformly distributed, as they can be scattered in reality.
@@ III-B2 Multidimensional scattered interpolation with RBFs
Although multiple approaches exist to construct $\mathcal{I}^{\mathcal{\mathcal{D}}}$ from the dataset $\mathcal{D}$, we propose the use of RBFs, as they provide an efficient and robust solution for multidimensional scattered interpolation [19, 20]. The RBF-based interpolant that approximates $\mathcal{L}_{f}$ is given by
@@
where $\phi$ is a RBF, i.e., a continuous function $\phi:\mathbb{R}^{d}\to\mathbb{C}$ with $d=n_{\mathrm{x}}+n_{\mathrm{u}}+n_{\mathrm{\eta}}$ which is invariant under translation and rotation, $\alpha\in\mathbb{R}^{N\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$, and the interpolation condition in (14a) is imposed as
@@
where $\mathbb{I}_{a}^{b}\coloneq\{i\in\mathbb{Z}^{+}\mid a\leq i\leq b\}$ is an index set. Then, the coefficients $\alpha_{i}$ can be determined by formulating (16) as
$R\;\alpha=\gamma,\vskip-5.0pt$ (17a)
where
$[R]_{i,j}\coloneq\phi(z^{(i)}-z^{(j)}),\quad i,j\in\mathbb{I}_{1}^{N},\quad R\in\mathbb{R}^{N\times N},$ (17b)
$[\gamma]_{i}\coloneq\operatorname{vec}(M_{i}),\quad i\in\mathbb{I}_{1}^{N},\quad\gamma\in\mathbb{R}^{N\times{n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}}.$ (17c)
@@ III-B3 Extension of the RBF interpolant with a polynomial tail
When variations of the local system dynamics are described by polynomial relationships of total degree less than $m$, the RBF-based interpolant (15) can be improved by adding a polynomial tail [13]. This polynomial tail represents the basis of $d$-variate polynomials of total degree at most m1m-1 which guarantees polynomial precision. The polynomial tail, denoted by $\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}$, is defined as
$\vskip-2.0pt\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z)=\operatorname{mat}_{n_{\mathrm{x}}}(\sum_{j=1}^{Q}\beta_{j}q_{j}(z)),$ (18)
where $\beta\in\mathbb{R}^{Q\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ and $q_{1},\dots,q_{Q}$ is a basis for the $d$-variate polynomials of degree $\leq m-1$ with dimension Q=(m1+d)!/((m1)!d!)Q={(m-1+d)!}/{((m-1)!\;d!)}. In this case, the interpolant that approximates $\mathcal{L}_{f}$ becomes
$\vskip-2.0pt\mathcal{I}^{\mathcal{D}}(z)=\mathcal{I}_{\mathrm{RBF}}^{\mathcal{\mathcal{D}}}(z)+\mathcal{I}_{\mathrm{poly}}^{\mathcal{D}}(z),$ (19)
@@
$\vskip-2.0pt\mathcal{I}^{\mathcal{D}}(z^{(i)})=M_{i},\quad i\in\mathbb{I}_{1}^{N},\\$ (20)
@@ III-B4 Determining coefficients of the interpolant
$\vskip-2.0pt\begin{bmatrix}R&P\\ P^{\top}&0\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\begin{bmatrix}\gamma\\ 0\end{bmatrix}$, (22a)
@@
$[P]_{i,j}$ $\coloneq q_{j}(z^{(i)}),\quad i\in\mathbb{I}_{1}^{N},\ j\in\mathbb{I}_{1}^{Q},\ P\in\mathbb{R}^{N\times Q},$ (22b)
where $R$ and $\gamma$ are defined as in (17). Now, we can show the uniqueness of solutions of (22) on the basis of [13, Chapter 8]. To this end, we first introduce the notion of conditionally positive definite (CPD) functions [13, Definition 8.1].
@@
Then, using Definition 3, we have the following result that closely follows the treatment of [13, Theorem 8.21].
@@
Suppose that the function $\phi$ is CPD of order $m$ and the columns of $P$ are linearly independent, where $P$ is given in (22b). Then, (22) has a unique solution.
@@
Suppose that $\alpha^{\top}\ \beta^{\top}]^{\top}\in\mathbb{R}^{(N+Q)\times n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})}$ is a solution of the homogeneous equation system in (22), i.e., with $\gamma=0$. Then, we have
$R\alpha+P\beta$ $=0,$ (25)
@@
where the second equation means that $\alpha$ satisfies the side condition (21). Now, pre-multiplying the first equation by $\alpha^{\top}$, giving $\alpha^{\top}\!R\alpha+\alpha^{\top}\!P\beta=\alpha^{\top}\!R\alpha+\left(P^{\top}\!\alpha\right)^{\!\top}\!\!\beta=\alpha^{\top}\!R\alpha=0$. As $\phi$ is CPD of order $m$, we can conclude that $\alpha=0$ and thus $P\beta=0$. Then, since the columns of $P$ are linearly independent, we can conclude that $\beta=0$. Therefore, (25) has trivial solutions only, meaning that the left-hand-side matrix of (22a) is full rank, and thus (22) has a unique solution. ∎
By Theorem 4, if the RBF is CPD and the data points $z^{(i)}$ yield a full-rank P in (22b), then (22) has a unique solution, ensuring that for a given $\mathcal{D}$ there exists a unique interpolant of the form (19) that satisfies (14a). One example of a CPD RBF function is the Hardy multiquadric [13], which is CPD of (minimum) order $m=1$, given by
$\phi(z;c)=-\left(c^{2}+\|z\|_{2}^{2}\right)^{1/2},$ (26)
@@
Now, for a given data set $\mathcal{D}$, we can use (15) to approximate the linearization operator $\mathcal{L}_{f}$ within the admissible space. Additionally, if the nonlinear system (1) is expected to contain polynomial relationships in its dynamics, we can use (19) instead to guarantee polynomial precision. In either case, the construction of the interpolant $\mathcal{I}^{\mathcal{D}}$ only involves an unconstrained linear system of equations, which can be efficiently solved with standard linear least squares methods [21]. In the next section, we detail how to utilize the proposed interpolant $\mathcal{I}^{\mathcal{D}}$ in combination with the results from Theorem 1 to achieve our objective of constructing the surrogate model on-the-fly.
@@ III-C Finding the surrogate representation on-the-fly
[…]
@@
$\hat{F}(\hat{x},u,\eta)\hskip-1.0pt$ $=\hskip-3.5pt\int_{0}^{1}\hskip-3.5pt\mathcal{I}^{\mathcal{D}}\left(\lambda\hat{x},\lambda u,\,\eta\right)\,\mathop{}\!\mathrm{d}\lambda,$ (27b)
and the integral in (27b) is left unevaluated. This representation allows assembling the surrogate model on-the-fly, i.e., the surrogate model is only computed when needed by solving the definite integral in (27b), as summarized in Algorithm 2. This is a key advantage for real-world applications, since incorporating new system observations only requires the recomputation of $\mathcal{I}^{\mathcal{D}}$. To this end, it is necessary to invoke a numerical integration scheme to solve (27b) in a computationally efficient manner, as an analytic solution can be costly and often not computable.
@@
Input: Data set $\mathcal{D}$ of the form (4), RBF function $\phi(z;c)$, width parameter $c$, polynomial basis $q(z)$ of total degree $m$.
@@
2:Find $\alpha$ and $\beta$ by solving (22).
@@
1:Compute $\hat{F}(\hat{x}(t),u(t),\eta)$ in (27b) with the numerical integration scheme.
2:Compute $\xi\hat{x}(t)$ by (27a).
@@
In addition, the proposed approach guarantees that $\hat{f}$ converges to the true nonlinear function $f$ as the number of observations in $\mathcal{D}$ increases, leading to the following result.
@@
Given a nonlinear system $\Sigma$ defined by (1) where $f\in\mathcal{C}^{k}$ with $k\in\mathcal{Z}^{+}$, then for a set $\mathcal{D}_{N}=\{\Theta,\mathcal{M}\}$ given by (4) with $N\in\mathbb{Z}^{+}$ distinct linearizations and the interpolant $\mathcal{I}^{\mathcal{D}}$ given by (19) that approximates $\mathcal{L}_{f}$ according to (2), the surrogate $\hat{f}$ given by (27a), satisfies
@@
at any point $(z_{\ast})\in\mathcal{Z}$ as $N\to\infty$.
@@
The proof follows from the fact that $\mathcal{I}^{\mathcal{D}}$ satisfies the interpolation condition in (14a), therefore $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|<\epsilon$ at any point $z\in\mathcal{Z}_{\Theta}^{\mathrm{c}}$, where $\mathcal{Z}_{\Theta}^{\mathrm{c}}$ denotes the relative complement of $\mathcal{Z}$ w.r.t. $\Theta$, i.e., $\mathcal{Z}_{\Theta}^{\mathrm{c}}\coloneq\mathcal{Z}\setminus\Theta$. Then, as $N\to\infty$, $\mathcal{Z}_{\Theta}^{\mathrm{c}}\to 0$, where #\# denotes the cardinality of the set. Thus, $\|\mathcal{L}_{f}-\mathcal{I}^{\mathcal{D}}\|\to 0$, and consequently, $\|f(z_{\ast})-\hat{f}(z_{\ast})\|\to 0$ by Theorem 1. ∎
@@ III-D Optimization of the RBFs width parameter
The RBF-based interpolation depends on the width $c$ (see (26)), which influences the approximation quality of the interpolant $\mathcal{I}^{\mathcal{D}}$ and, consequently, the accuracy of the surrogate $\hat{f}$. To tune $c$, we use the “leave-one-out” cross-validation based optimization from [14]. For a given $\mathcal{D}_{N}=\{z^{(k)},M_{k}\}_{k=1}^{N}$, an RBF $\phi(z;c)$ and a polynomial basis $q$ of total degree $m$,
$\vskip-5.0ptc^{\ast}=\arg\min_{c}\|[E_{1}\ldots E_{N}](c)\|_{p},$ (29)
@@
where $I^{\mathcal{D}_{k}}$ denotes the interpolant given by (19) over the set $\mathcal{D}_{k}\coloneq\mathcal{D}\setminus\{z^{(k)}\}$, and $\|\cdot\|_{p}$ denotes the induced $p$ matrix norm. In particular, $p=\infty$ corresponds to (14b). Note that (29) is nonlinear due to the norm and the matrix inversion required for $I^{\mathcal{D}_{k}}$, and non-smooth when $p=\infty$.